OpenStudy (anonymous):

solve and write using set builder notation. 12-8(3x-5)>-7x+29

OpenStudy (paxpolaris):

\[\Large12-8\left( 3x-5\right)>-7x+29\] multiply out: -8(3x-5) \[-8(3x-5) =-24x+40\] \[\implies\Large12-24x+40>-7x+29\]\[\implies\Large-24x+52>-7x+29\] Next, bring all the x-terms on one side and all the no-x-terms to the other side: \[\implies\Large52-29>-7x+24x\]\[\implies\Large23>17x\] Lastly divide both sides by 17\[\huge \therefore{23\over17}>x\] or \[\huge x<{23\over17}\] in set notation: \[\huge \left\{ x \in \mathbb{R}:x<{23 \over7} \right\}\]

OpenStudy (anonymous):

Ok, Thank You wanted to check my answer.